OpenStudy (anonymous):

Write an explicit formula for the sequence 1/2, 3/7, 1/3, 5/19, 3/14... then find a14+ Can anyone help me? Im so lost

OpenStudy (anonymous):

@JamesJ @satellite73 @Campbell_st

OpenStudy (anonymous):

I will give out a medal

OpenStudy (anonymous):

This problem can have one solution like \[ a_n=-\frac{n^4}{456}+\frac{4 n^3}{133}-\frac{439 n^2}{3192}+\frac{87 n}{532}+\frac{178}{399} \] I am sure whoever asked you this question does not expect this answer. But this answer is correct

OpenStudy (anonymous):

\[ a_{14}= -\frac{3450}{133} \]

OpenStudy (anonymous):

I am sure this problem has many solutions.

OpenStudy (anonymous):

My choices are @eliassaab an=an-1 - (n-1)/7n; 15/199 an= (an=1)/(n^2+3); 15/199 an= (n+1)/(n^2+3); 15/199 an= n/n^3-1; 15/199

OpenStudy (anonymous):

The third choice

OpenStudy (anonymous):

Try it for n=1, you find 1/2

OpenStudy (anonymous):

Thank you so much!!!

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

Did you try n=2,3,4,5 and 14

OpenStudy (anonymous):

C